OpenStudy (grant330sims):

describe how you would verify the zeros of the function x3 – 3x2 – 16x + 48?

sam (.sam.):

I would use a calculator to find a zero first

OpenStudy (grant330sims):

I know what the zero's are, its just that i dont understand what it means by verify

sam (.sam.):

oh just expand your factored term

OpenStudy (grant330sims):

just listing the factors with the + or -?

sam (.sam.):

Lets say the zeros are x=3 x=4 x=-4 then (x-3)(x-4)(x+4) then expand and verify if you got the same equation x3 – 3x2 – 16x + 48

OpenStudy (mertsj):

\[x^3-3x^2-16x+48=x^2(x-3)-16(x-3)=(x-3)(x^2-16)=(x-3)(x-4)(x+4)=0\]

OpenStudy (mertsj):

\[(x-3)(x^2-16)=(x-3)(x-4)(x+4)=0\]

OpenStudy (mertsj):

If the product of factors is zero, one of the factors is 0 so: x-3=0 or x-4=0 or x+4=0 Solving, we have: x-3, x=4, x=-4 We have just verified the zeros of the function.

OpenStudy (grant330sims):

Thanks man!

OpenStudy (mertsj):

yw